Vapour density of N2O4 at 60 degree centigrade is found to be 306. Find out degree of dissociation of N2O4.

Grade:12

1 Answers

vikas askiitian expert
13 years ago

vapour dencity = 2(molar mass)

                 dc(calculated)   =2(28+64)=184

de = degree of dissociation experimental

dc/de = degree of dissotiation

(184/306)*100 = degree of dissociation